OpenStudy (anonymous):

A social security number has seven digits. How many Social Security numbers are possible?

jimthompson5910 (jim_thompson5910):

Ideally there are 10^7 = 10,000,000 (ten million) possible social security numbers. However, I'm betting there are large chunk of numbers that aren't allowed (like ones that start with 000 maybe?), but I'm not sure which ones.

OpenStudy (unicornpoopcookies):

Suppose a social security number is ABC-DE-FGHI There are 10 choices for a digit to go in the A position There are 10 choices for a digit to go in the B position There are 10 choices for a digit to go in the C position There are 10 choices for a digit to go in the D position There are 10 choices for a digit to go in the E position There are 10 choices for a digit to go in the F position There are 10 choices for a digit to go in the G position There are 10 choices for a digit to go in the H position There are 10 choices for a digit to go in the I position So that's 10*10*10*10*10*10*10*10*10 = 1000000000. *check link*

jimthompson5910 (jim_thompson5910):

hmm I guess the problem liltemac32 posted is a bit off, but it does say 7 digits, not 9